the areA of square plate is calculated by measuring the length to be 5.5 cm. if the scale used has the leAST COUNT OF 0.1 CM THEN WHAT IS THE MAXIMUM PERCENTAGE ERROR IN THE AREA?
 
 

 

Grade:11

1 Answers

Shailendra Kumar Sharma
6 years ago
Area of reactangle is l^2 where l is side
Δl/l+ Δl/l = ΔA/A 
Now .1/5.5 +.1/5.5=ΔA/A  =.0366
So % error  in Area will be 3.66%